# A Culturally-diverse Multilingual Multimodal Video Benchmark & Model

### v1

@@ Abstract
Large multimodal models (LMMs) have recently gained attention due to their effectiveness to understand and generate descriptions of visual content. Most existing LMMs are in English language. While few recent works explore multilingual image LMMs, to the best of our knowledge, moving beyond the English language for cultural and linguistic inclusivity is yet to be investigated in the context of video LMMs. In pursuit of more inclusive video LMMs, we introduce a multilingual Video LMM benchmark, named ViMUL-Bench, to evaluate Video LMMs across 14 languages, including both low- and high-resource languages: English, Chinese, Spanish, French, German, Hindi, Arabic, Russian, Bengali, Urdu, Sinhala, Tamil, Swedish, and Japanese. Our ViMUL-Bench is designed to rigorously test video LMMs across 15 categories including eight culturally diverse categories, ranging from lifestyles and festivals to foods and rituals and from local landmarks to prominent cultural personalities. ViMUL-Bench comprises both open-ended (short and long-form) and multiple-choice questions spanning various video durations (short, medium, and long) with 8k samples that are manually verified by native language speakers. In addition, we also introduce a machine translated multilingual video training set comprising 1.2 million samples and develop a simple multilingual video LMM, named ViMUL, that is shown to provide a better tradeoff between high-and low-resource languages for video understanding. We hope our ViMUL-Bench and multilingual video LMM along with a large-scale multilingual video training set will help ease future research in developing cultural and linguistic inclusive multilingual video LMMs. Our proposed benchmark, video LMM and training data are publicly released at https://mbzuai-oryx.github.io/ViMUL/
@@ A Culturally-diverse Multilingual Multimodal Video Benchmark & Model
Bhuiyan Sanjid Shafique1 Ashmal Vayani1,2, Muhammad Maaz1, Hanoona Abdul Rasheed1, Dinura Dissanayake1, Mohammed Irfan Kurpath1, Yahya Hmaiti2, Go Inoue1, Jean Lahoud1, Md. Safirur Rashid3, Shadid Intisar Quasem3, Maheen Fatima4, Franco Vidal2, Mykola Maslych2, Ketan Pravin More1, Sanoojan Baliah1, Hasindri Watawana1, Yuhao Li1, Fabian Farestam5, Leon Schaller6, Roman Tymtsiv7, Simon Weber6, Hisham Cholakkal1, Ivan Laptev1, Shin’ichi Satoh8, Michael Felsberg10, Mubarak Shah2,        Salman Khan1,9, Fahad Shahbaz Khan1,10 1Mohamed bin Zayed University of AI, 2University of Central Florida, 3Islamic University of Technology, 4Air University, 5ETH Zurich, 6 Technische Universität München, 7Independent Researcher, 8National Institute of Informatics, 9Australian National University, 10Linköping University
@@ 1 Introduction
Large Multimodal Models (LMMs) have achieved remarkable success in vision-and-language tasks, yet their development has predominantly centered on English, overlooking the vast linguistic and cultural diversity of global users Vayani et al. (2025); Pfeiffer et al. (2021). This English-centric focus leads to significant performance gaps for other languages, as models often fail to grasp user intent when queries or captions are in low-resource languages. Moreover, current models struggle with cultural nuances and region-specific context specifically for low-resource languages Romero et al. (2024). For instance, the MaRVL Liu et al. and the recent ALM-Bench Vayani et al. (2025) image LMM benchmarks, verified by native speakers to include diverse (including low-resource) languages, reveal dramatic drops in accuracy when state-of-the-art models operate beyond English. These findings underscore the pressing need for multilingual and multicultural evaluation benchmarks to develop more inclusive next generation of LMMs.
Existing efforts to explore linguistically and culturally diverse LMM benchmarks are limited to images Romero et al. (2024); Vayani et al. (2025). To the best of our knowledge, linguistic and cultural diversity are yet to be investigated for video LMMs. Video domain poses different challenges as it often depicts complex, culturally rich scenarios—local festivals, foods, rituals, or landmarks, that require understanding both the visual context and the language-specific narration or questions. Fig. 1 illustrates an example where LMMs are asked to describe the taste of the Bengali dish Beef Tehari. The models fail to interpret the question in the local language, responding incorrectly and missing linguistic nuances. While short and long video understanding LMM benchmarks exist in literature, they are typically restricted to only English language. For instance, Video-MME Fu et al. (2024) focuses on diverse video analysis but in a single language, and MVBench Li et al. (2024b) emphasizes temporal reasoning (action sequences, motion) without multilingual considerations. Other recent efforts like ViLMA Kesen et al. (2024) and SEED-Bench Li et al. (2023b, a) probe video-language models’ abilities in zero-shot temporal grounding and procedural understanding, among other skills, yet none assess cross-lingual or cross-cultural comprehension. In short, there is no comprehensive benchmark to evaluate how well video LMMs perform across different languages and cultural contexts (see Tab. 1).
@@ 2 Related Works
ActivityNet-QA Fabian Caba Heilbron and Niebles (2015) (1) - 2378 OE 5800 Human
@@
Table 1: Comparison of video LMM benchmarks emphasizing multilingual and cultural understanding. Domains represent the aspects covered by each dataset for different languages. Annotation Type is categorized as follows: Human - Questions were created in the local language. Human+Auto – Questions were generated or translated using GPT-4/Google API and later validated by human experts. Auto: Questions were generated or translated automatically without human validation. ‘-’ indicates that information is not available.
@@
Video LMM Benchmarks: Extending multimodal evaluation to video introduces additional challenges due to temporal dynamics. A number of video-language benchmarks have emerged to assess LMMs on video understanding Mahmood et al. (2024), though they focus on general monolingual capabilities. Video-MME Fu et al. (2024) is introduced as the first comprehensive evaluation of LMMs on video analysis, covering a broad spectrum of video domains. It offers a full-spectrum evaluation, covering questions from perception to reasoning, and variable videos lengths ($\sim$10 secs to $\sim$1 hour). Another effort, MVBench Li et al. (2024b) concentrates on temporal reasoning skills and proposes an automatic pipeline to generate a large multiple-choice QA benchmark by leveraging existing video datasets and GPT-based annotators. Beyond general benchmarks, ViLMA Kesen et al. (2024) takes a fine-grained approach as it uses carefully designed counterfactual video pairs to probe a model’s temporal grounding and linguistic understanding in a zero-shot setting. SEED-Bench is another comprehensive multimodal benchmark that includes some video-based questions; however, those are largely confined to temporal or procedural understanding tasks Li et al. (2023b, a). These benchmarks have a significantly advanced evaluation for video-based multimodal reasoning and generation, however they lack multilingual and culturally aware components. Current video benchmarks evaluate models primarily on English video-question pairs and generic content, without testing performance on non-English dialogues, region-specific contexts, or culturally diverse visual narratives. The absence of multilingual and multicultural evaluation for video LMMs forms a key motivation for ViMUL-Bench, which aims to fill that gap by assessing video understanding across diverse languages and cultural settings.
@@ 3 ViMUL-Bench
ViMUL-Bench is a comprehensive multilingual benchmark, designed to evaluate both general and culturally-specific aspects of video comprehension in video LMMs. It captures rich cultural nuances through a diverse set of question-answer (QA) pairs, including multiple-choice and open-ended (short and long) formats. The benchmark spans 15 diverse categories, categorized into general and cultural topics, across 14 languages: Arabic, Bengali, Chinese, English, French, German, Hindi, Japanese, Russian, Sinhala, Spanish, Swedish, Tamil, and Urdu. We build upon the recent PALO Maaz et al. (2025), incorporating its 10 languages while adding Swedish, German, Tamil, and Sinhala to ensure typological diversity and to enhance the representation of low-resource languages such as Tamil, Urdu, and Sinhala as defined by Ethnologue Campbell and Grondona (2008) and Glottolog Hammarström et al. (2022) database.
The generic category includes seven domains: Artistic Performance, Digital Content, Knowledge, Life Record, Sports Competition, Surveillance, and Others. The cultural categorization is inspired from recent image LMM benchmarks Vayani et al. (2025); Romero et al. (2024); Marino et al. (2019), where the corresponding videos for each domain and language are manually scrapped with their annotations manually verified by a native speaker. It spans eight diverse domains, including Lifestyle & Customs, Festivals, Food & Cuisine, Sports, Architecture & Landmarks, Notable Public Figures, Media & Entertainment, and Art & Literature. Our ViMUL-Bench has been meticulously curated and verified by native experts for the 13 languages to ensure high-quality question answer (QA) pairs that accurately capture the nuances of all 15 domains Campos et al. (2025); Raza et al. (2025a). It comprises 8,025 diverse questions in total across 14 languages, spanning both generic and cultural categories to comprehensively evaluate multilingual and cross-cultural video understanding.
@@ 3.1 Data Collection and Annotation
Cultural Video Curation. To curate diverse cultural QA pairs, we collect open-licensed videos and their corresponding metadata from the internet Raza et al. (2025c), focusing on specific cultural aspects of each language across three durations: short (0-4 mins), medium (4-15 mins), and long (15+ mins). To accurately capture the cultural nuances of each language, we follow the approach outlined in Vayani et al. (2025); Romero et al. (2024) and link each language to a country based on the World Values Survey Haerpfer et al. (2022) to ensure coverage of cultural and ritual diversity. Additionally, we generate a list of topics for each category, forming a triplet (language-country-topic), using GPT-4o, and then search for relevant content. For example, querying “Popular sports in United States" yields responses such as “baseball, soccer, golf, and ice hockey." We manually extract videos in the native language to ensure linguistic accuracy. Each domain undergoes several filtration steps, such as removing low-resolution, noisy, or unclear videos, to guarantee data quality. To maintain both high-quality and cultural relevance, we enlist expert native speakers of each language to manually verify the quality and cultural diversity of videos. Any content lacking cultural relevance is removed from the dataset. Fig. 3 shows our data collection and verification pipeline.
Cultural QA Generation. To generate high-quality video-QA pairs for the cultural section of ViMUL-Bench, the question-answer (QA) pairs are curated via native experts based on the provided videos and their metadata. Notably, videos and their corresponding QAs are not shared across languages in the cultural set. For each video, we generate multiple-choice questions (MCQs) and one open-ended question in the native language. To reduce randomness in MCQ generation, we ensure that each question can also be answered when rephrased as an open-ended question. Further, the native experts are tasked with creating an English version of the question-answer pair Raza et al. (2025b). We instruct the experts to focus on cultural concepts depicted in each video and to generate questions that require a visual understanding of the video, while avoiding the perpetuation of bias and stereotypes. Following this process, 2,633 QA pairs were curated, spanning 337 videos for the cultural category.
@@ 4.1 Multilingual Instruction Tuning Dataset
Figure 6: We present qualitative examples of failure cases of GPT-4o’s across different language scripts and categories, specifying the corresponding error types. For instance, in a Sinhala-language question asking about the event being celebrated in the video, the model correctly identifies the cultural significance—celebrating the birth, enlightenment, and passing of Lord Buddha, but fails to respond with grammatically correct Sinhala , highlighting a language proficiency error. Results on success cases are shown in Fig. 11 (suppl. material).
@@ 5 Results and Discussions
As discussed earlier, ViMUL-Bench comprises two question types: multiple-choice and open-ended, including both short and long question-answer formats, where we employ different prompts for each question type. For multiple-choice questions, we provide the visual context and textual query to the LMMs, instructing them to select the best option, which is then directly compared to the ground truth. Performance is measured using accuracy, following established multiple-choice VQA benchmarks Romero et al. (2024); Bang et al. (2023); Zhu et al. (2016). For open-ended questions, we use the open-source multilingual LLM, Phi-4-14B Abdin et al. (2024) as a judge, ensuring consistency and reproducibility, unlike GPT-based models Shen et al. (2023); Stureborg et al. (2024), which are costly and inconsistent due to version updates. Performance is evaluated using correctness criteria, which measure how closely the model’s output matches the ground truth Vayani et al. (2025); Narnaware et al. (2025); Thawakar et al. (2024).
@@
Performance across Language Scripts. We group the 14 languages in our ViMUL-Bench by language scripts Costa-Jussà et al. (2022), using data from Ethnologue Campbell and Grondona (2008) and the Glottolog database Hammarström et al. (2022). This results in nine distinct scripts. Fig. 8 shows the performance of video LMMs across these scripts. The closed-source GPT-4o consistently achieves the best results across all scripts. Among open-source models, ViMUL performs favorably against all existing methods across all language scripts, with accuracy ranging from  32% to  55%. Additionally, all video LMMs struggle with Sinhalese (Sinhala) and Tamil scripts, with ViMUL outperforming the second-best open-source model by  8%. A performance gradient is observed, with video LMM performing significantly better on Latin, Chinese, and Cyrillic scripts compared to Sinhalese (Sinhala) and Tamil.

### v2

@@ Abstract
Large multimodal models (LMMs) have recently gained attention due to their effectiveness to understand and generate descriptions of visual content. Most existing LMMs are in English language. While few recent works explore multilingual image LMMs, to the best of our knowledge, moving beyond the English language for cultural and linguistic inclusivity is yet to be investigated in the context of video LMMs. In pursuit of more inclusive video LMMs, we introduce a multilingual Video LMM benchmark, named ViMUL-Bench, to evaluate Video LMMs across 14 languages, including both low- and high-resource languages: English, Chinese, Spanish, French, German, Hindi, Arabic, Russian, Bengali, Urdu, Sinhala, Tamil, Swedish, and Japanese. Our ViMUL-Bench is designed to rigorously test video LMMs across 15 categories including eight culturally diverse categories, ranging from lifestyles and festivals to foods and rituals and from local landmarks to prominent cultural personalities. ViMUL-Bench comprises both open-ended (short and long-form) and multiple-choice questions spanning various video durations (short, medium, and long) with 8k samples that are manually verified by native language speakers. In addition, we also introduce a machine translated multilingual video training set comprising 1.2 million samples and develop a simple multilingual video LMM, named ViMUL, that is shown to provide a better tradeoff between high-and low-resource languages for video understanding. We hope our ViMUL-Bench and multilingual video LMM along with a large-scale multilingual video training set will help ease future research in developing cultural and linguistic inclusive multilingual video LMMs. Our proposed benchmark, video LMM codebase and training data are available at https://mbzuai-oryx.github.io/ViMUL/.
@@ A Culturally-diverse Multilingual Multimodal Video Benchmark & Model
Bhuiyan Sanjid Shafique1thanks: Equal contribution, Ashmal Vayani211footnotemark: 1, Muhammad Maaz1, Hanoona Abdul Rasheed1, Dinura Dissanayake1, Mohammed Irfan Kurpath1, Yahya Hmaiti2 Go Inoue1, Jean Lahoud1, Md. Safirur Rashid3, Shadid Intisar Quasem3, Maheen Fatima4 Franco Vidal2, Mykola Maslych2, Ketan Pravin More1, Sanoojan Baliah1 Hasindri Watawana1, Yuhao Li1, Fabian Farestam5, Leon Schaller6, Roman Tymtsiv7 Simon Weber6, Hisham Cholakkal1, Ivan Laptev1, Shin’ichi Satoh8, Michael Felsberg10 Mubarak Shah2, Salman Khan1,9, Fahad Shahbaz Khan1,10 1Mohamed bin Zayed University of AI, 2University of Central Florida, 3Islamic University of Technology, 4Air University 5ETH Zurich, 6 Technische Universität München, 7Independent Researcher, 8National Institute of Informatics 9Australian National University, 10Linköping University {bhuiyan.shafique, muhammad.maaz, hanoona.bangalath, dinura.dissanayake, mohammedirfan.k, go.inoue, jean.lahoud ketan.more, sanoojan.baliah, hasindri.watawana, hisham.cholakkal, salman.khan, fahad.khan}@mbzuai.ac.ae {ashmal.vayani, yohan.hmaiti, fr543419, mykola.maslych}@ucf.edu, {safirurrashid, shadidintisar}@iut-dhaka.edu {leon.schaller, simon1.weber}@tum.de, {tymtsiv.roman, lyh88524}@gmail.com, maheen.fatima@students.au.edu.pk ffarestam@student.ethz.ch, satoh@nii.ac.jp, michael.felsberg@liu.se, ivan.laptev@inria.fr, mubarak.shah@crcv.ucf.edu
@@ 1 Introduction
Large Multimodal Models (LMMs) have achieved remarkable success in vision-and-language tasks, yet their development has predominantly centered on English, overlooking the vast linguistic and cultural diversity of global users Vayani et al. (2025); Pfeiffer et al. (2021). This English-centric focus leads to significant performance gaps for other languages, as models often fail to grasp user intent when queries or captions are in low-resource languages. Moreover, current models struggle with cultural nuances and region-specific context specifically for low-resource languages Romero et al. (2024); Qureshi et al. (2025); Raza et al. (2025a). For instance, the MaRVL Liu et al. and the recent ALM-Bench Vayani et al. (2025) image LMM benchmarks, verified by native speakers to include diverse (including low-resource) languages, reveal dramatic drops in accuracy when state-of-the-art models operate beyond English. These findings underscore the pressing need for multilingual and multicultural evaluation benchmarks to develop more inclusive next generation of LMMs.
Existing efforts to explore linguistically and culturally diverse LMM benchmarks are limited to images Romero et al. (2024); Vayani et al. (2025). To the best of our knowledge, linguistic and cultural diversity are yet to be investigated for video LMMs. Video domain poses different challenges as it often depicts complex, culturally rich scenarios-local festivals, foods, rituals, or landmarks, that require understanding both the visual context and the language-specific narration or questions. Fig. 1 illustrates an example where LMMs are asked to describe the taste of the Bengali dish Beef Tehari. The models fail to interpret the question in the local language, responding incorrectly and missing linguistic nuances. While short and long video understanding LMM benchmarks exist in literature, they are typically restricted to only English language. For instance, Video-MME Fu et al. (2024) focuses on diverse video analysis but in a single language, and MVBench Li et al. (2024b) emphasizes temporal reasoning (action sequences, motion) without multilingual considerations. Other recent efforts like ViLMA Kesen et al. (2024) and SEED-Bench Li et al. (2023b, a) probe video-language model’s abilities in zero-shot temporal grounding and procedural understanding, among other skills, yet none assess cross-lingual or cross-cultural comprehension. In short, there is no comprehensive benchmark to evaluate how well video LMMs perform across different languages and cultural contexts (see Tab. 1).
@@ 2 Related Works
ActivityNet-QA
Fabian Caba Heilbron and Niebles (2015) (1) - 2378 OE 5800 Human
@@
Table 1: Comparison of video LMM benchmarks emphasizing multilingual and cultural understanding. Domains represent the aspects covered by each dataset for different languages. Annotation Type is categorized as follows: Human - Questions were created in the local language. Human+Auto - Questions were generated or translated using GPT-4/Google API and later validated by human experts. Auto: Questions were generated or translated automatically without human validation. ’-’ indicates that information is not available.
@@
Video LMM Benchmarks: Extending multimodal evaluation to video introduces additional challenges due to temporal dynamics. A number of video-language benchmarks have emerged to assess LMMs on video understanding, though they focus on general monolingual capabilities. Video-MME Fu et al. (2024) is introduced as the first comprehensive evaluation of LMMs on video analysis, covering a broad spectrum of video domains. It offers a full-spectrum evaluation, covering questions from perception to reasoning, and variable videos lengths ($\sim$10 secs to $\sim$1 hour). Another effort, MVBench Li et al. (2024b) concentrates on temporal reasoning skills and proposes an automatic pipeline to generate a large multiple-choice QA benchmark by leveraging existing video datasets and GPT-based annotators. Beyond general benchmarks, ViLMA Kesen et al. (2024) takes a fine-grained approach as it uses carefully designed counterfactual video pairs to probe a model’s temporal grounding and linguistic understanding in a zero-shot setting. SEED-Bench is another comprehensive multimodal benchmark that includes some video-based questions; however, those are largely confined to temporal or procedural understanding tasks Li et al. (2023b, a). These benchmarks have a significantly advanced evaluation for video-based multimodal reasoning and generation, however they lack multilingual and culturally aware components. Current video benchmarks evaluate models primarily on English video-question pairs and generic content, without testing performance on non-English dialogues, region-specific contexts, or culturally diverse visual narratives. The absence of multilingual and multicultural evaluation for video LMMs forms a key motivation for ViMUL-Bench, which aims to fill that gap by assessing video understanding across diverse languages and cultural settings.
@@ 3 ViMUL-Bench
ViMUL-Bench is a comprehensive multilingual benchmark, designed to evaluate both general and culturally-specific aspects of video comprehension in video LMMs. It captures rich cultural nuances through a diverse set of question-answer (QA) pairs, including multiple-choice and open-ended (short and long) formats Raza et al. (2025b); Campos et al. (2025). The benchmark spans 15 diverse categories, categorized into general and cultural topics, across 14 languages: Arabic, Bengali, Chinese, English, French, German, Hindi, Japanese, Russian, Sinhala, Spanish, Swedish, Tamil, and Urdu. We build upon the recent PALO Maaz et al. (2025), incorporating its 10 languages while adding Swedish, German, Tamil, and Sinhala to ensure typological diversity and to enhance the representation of low-resource languages such as Tamil, Urdu, and Sinhala as defined by Ethnologue Campbell and Grondona (2008) and Glottolog Hammarström et al. (2022) database.
The generic category includes seven domains: Artistic Performance, Digital Content, Knowledge, Life Record, Sports Competition, Surveillance, and Others. The cultural categorization is inspired from recent image LMM benchmarks Vayani et al. (2025); Romero et al. (2024); Marino et al. (2019), where the corresponding videos for each domain and language are manually scrapped with their annotations manually verified by a native speaker. It spans eight diverse domains, including Lifestyle & Customs, Festivals, Food & Cuisine, Sports, Architecture & Landmarks, Notable Public Figures, Media & Entertainment, and Art & Literature. Our ViMUL-Bench has been meticulously curated and verified by native experts for the 13 languages to ensure high-quality question answer (QA) pairs that accurately capture the nuances of all 15 domains. It comprises 8,025 diverse questions in total across 14 languages, spanning both generic and cultural categories to comprehensively evaluate multilingual and cross-cultural video understanding.
@@ 3.1 Data Collection and Annotation
Cultural Video Curation. To curate diverse cultural QA pairs, we collect open-licensed videos and their corresponding metadata from the internet, focusing on specific cultural aspects of each language across three durations: short (0-4 mins), medium (4-15 mins), and long (15+ mins). To accurately capture the cultural nuances of each language, we follow the approach outlined in Vayani et al. (2025); Romero et al. (2024) and link each language to a country based on the World Values Survey Haerpfer et al. (2022) to ensure coverage of cultural and ritual diversity. Additionally, we generate a list of topics for each category, forming a triplet (language-country-topic), using GPT-4o, and then search for relevant content. For example, querying “Popular sports in United States" yields responses such as “baseball, soccer, golf, and ice hockey." We manually extract videos in the native language to ensure linguistic accuracy. Each domain undergoes several filtration steps, such as removing low-resolution, noisy, or unclear videos, to guarantee data quality. To maintain both high-quality and cultural relevance, we enlist expert native speakers of each language to manually verify the quality and cultural diversity of videos. Any content lacking cultural relevance is removed from the dataset. Fig. 3 shows our data collection and verification pipeline.
Cultural QA Generation. To generate high-quality video-QA pairs for the cultural section of ViMUL-Bench, the question-answer (QA) pairs are curated via native experts based on the provided videos and their metadata. Notably, videos and their corresponding QAs are not shared across languages in the cultural set. For each video, we generate multiple-choice questions (MCQs) and one open-ended question in the native language. To reduce randomness in MCQ generation, we ensure that each question can also be answered when rephrased as an open-ended question. Further, the native experts are tasked with creating an English version of the question-answer pair. We instruct the experts to focus on cultural concepts depicted in each video and to generate questions that require a visual understanding of the video, while avoiding the perpetuation of bias and stereotypes. Following this process, 2,633 QA pairs were curated, spanning 337 videos for the cultural category.
@@ 4.1 Multilingual Instruction Tuning Dataset
Figure 6: We present qualitative examples of failure cases of GPT-4o’s across different language scripts and categories, specifying the corresponding error types. For instance, in a Sinhala-language question asking about the event being celebrated in the video, the model correctly identifies the cultural significance-celebrating the birth, enlightenment, and passing of Lord Buddha, but fails to respond with grammatically correct Sinhala, highlighting a language proficiency error. Results on success cases are shown in Fig. 9 (suppl. material).
@@ 5 Results and Discussions
As discussed earlier, ViMUL-Bench comprises two question types: multiple-choice and open-ended, including both short and long question-answer formats, different prompts are employed for each question type. For multiple-choice questions, we provide the visual context and textual query to the LMMs, instructing them to select the best option, which is then directly compared to the ground truth. Performance is measured using accuracy, following established multiple-choice VQA benchmarks Romero et al. (2024); Bang et al. (2023); Zhu et al. (2016). For open-ended questions, we use the open-source multilingual LLM, Phi-4-14B Abdin et al. (2024) as a judge, ensuring consistency and reproducibility, unlike GPT-based models Shen et al. (2023); Stureborg et al. (2024), which are costly and inconsistent due to version updates. Performance is evaluated using correctness criteria, which measure how closely the model’s output matches the ground truth (see Sec. A in suppl. material for further detail). This approach follows recent work in evaluation frameworks Vayani et al. (2025); Narnaware et al. (2025), though the correctness used here is specific to our setup.
@@
Performance across Language Scripts. We group the 14 languages in our ViMUL-Bench by language scripts Costa-Jussà et al. (2022), using data from Ethnologue Campbell and Grondona (2008) and the Glottolog database Hammarström et al. (2022). This results in nine distinct scripts. Fig. 8 shows the performance of video LMMs across these scripts. The closed-source GPT-4o consistently achieves the best results across all scripts. Among open-source models, ViMUL performs favorably against all existing methods across all language scripts, except Chinese, with accuracy ranging from  32% to  55%. Additionally, all video LMMs struggle with Sinhalese (Sinhala) and Tamil scripts, with ViMUL outperforming the second-best open-source model by  8%. A performance gradient is observed, with video LMM performing significantly better on Latin, Chinese, and Cyrillic scripts compared to Sinhalese (Sinhala) and Tamil.